During the cleanup of the lab. Steve found a box with two BNCs going to the ICS DAQ interface and an unconnected D-SUB on the floor under the AP table. It seemed like a temperature sensor.
The BNCs were connected to C1:PEM-OSA_APTEMP and C1:PEM-OSA_SPTEMP.
Steve removed the box from the floor. These channels can be now used as spare DAQ channels. I labeled those cables.